OpenStudy (anonymous):

. Terry is mailing two boxes. Together they weigh 21 lbs. If the smaller box is 5 lbs. less than the larger one, how much does each box weigh? Write and solve an equation that models this scenario.

OpenStudy (anonymous):

Let small box = x Let large box = y Both boxes together weigh 21 lbs. In other words, x + y = 21 Small box is 5 lbs less than large box. In other words, x + 5 = y Now isolate a variable and you can solve

OpenStudy (bahrom7893):

if y=x+5 how can you represent x+y=21?

OpenStudy (bahrom7893):